------------------------------------------------------------ Date: Wed Oct 31 08:17:56 2001, Shift: 00:00-08:00 Supervisor: DO, with: Log book#10 pages 76 - 78 Shift Summary: ------------------------------------------------------------ 1.Machine status The good store started at 21:45 continued until 1:30 The beam was dumped without warning. Around 2:40, there was a new injection and ramping. MCR could store it. But the ZDC coincidence rate / ZDC single was about 1/50 with a low intensity (started around 150Hz and decreased around 50). All experiments had this low ZDC rate. I didn't take any data from this store because of the background. MCR called to tell they had a problem with blue beam life time. A problem occured in the AGS main magnet cooling. They needed 1 hour to fix all that. A 4:51, new injection and ramping. The beam survived at reasonable intensities (B28, Y20) with a ZDC coincidence rate ~ 920 Hz (ratio with ZDC single ~ 1/15). MCR called to tell they wanted to rebucket the beam. Beam survived after rebucketing. They finally dumped the beam at 5:50 for beam studies that will continue until 18:00 today. 2.Achievements: Runs Trig6 Spectrometer Setting 5593 70k FS 8deg 1/5 A, MRS 45deg 4B 5594 89k same setting New setting: FS @ 12deg, 1/4 A, MRS @ 45deg 4A TPC pedestal run 5596, DVM 5597 Note: I wanted to set 1/8 field in FS but couldn't find the current values. Did not think that I should just divide the 1/4 setting by 2, so skipped the 1/8 setting for the 1/4 program. 3.No detector problems this night! TOFW TDC top 92 still absent. T5 mod II off. ------------------------------------------------------------
This archive was generated by hypermail 2b30 : Wed Oct 31 2001 - 08:18:53 EST